OpenStudy (anonymous):

FAN AND MEDAL Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. Part B: Use the formula for area and solve for the value of pi for each circle. Part C: What observation can you make about the value of pi for circles A and B?

OpenStudy (anonymous):

@starlord6200

OpenStudy (anonymous):

@william.1233 can u help me?

OpenStudy (anonymous):

wow this one is hard

OpenStudy (anonymous):

c=2pi *r=25.12 ....pi=25.12/2*4 c=2pi*r=9.42 ....pi=9.42/2*1.5